OpenStudy (anonymous):

2(11) IS THIS COMMUTATIVE OR DISTRIBUTIVE PROPERTY

OpenStudy (anonymous):

commutative property !!! I think

OpenStudy (anonymous):

WHAT ABOUT 2(11+4)

OpenStudy (anonymous):

yes thats distributive 2*11 + 2*4

OpenStudy (anonymous):

WHAT ABOUT (M+4)-5=3+(4-5) AND 2N+P=P+2N

OpenStudy (anonymous):

associative property

OpenStudy (anonymous):

FOR WHICH ONE

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

WHAT ABOUT THE 2ND ONE

OpenStudy (anonymous):

commutative

OpenStudy (anonymous):

SIMPLIFY EXPRESSION. USE ONLY POSITIVE EXPONENTS \[4^{7} OVER 4^{5}\]

OpenStudy (anonymous):

4^2= 16

OpenStudy (anonymous):

what about 12m to the 6thpower over 5m to the 6thpower